OpenStudy (anonymous):

What are the coordinates of the x-intercepts of the graph of y = 2x2 + 6x – 20? HELLPP! I am so incredibly confused!

OpenStudy (anonymous):

y = 2x^2 + 6x - 20 -- set y equal to zero 0 = 2x^2 + 6x - 20 = 2 (x^2 + 6x - 20) --- 2 cannot equal zero so.. x^2 + 6x - 20 = 0 ( x - 2 ) ( x + 5 ) = 0 x - 2 = 0 --> x=2 x + 5 = 0 --> x=-5 plug into original equation to find the y intercepts x=2: y = 2(2)^2 + 6(2) -20 = 0 x=-5: y = 2(-5)^2 + 6(-5) -20 = 0 coordinates for x-intercepts (2,0) and (-5,0)

OpenStudy (anonymous):

Thank you so much this is perfect Cx